OpenStudy (anonymous):

Solve: negative 1 over 2 x + 3 = –x + 7

OpenStudy (darkigloo):

\[\frac{ -1 }{ 2x+3 } = -x + 7\] multiply both sides by (2x+3)

OpenStudy (anonymous):

ok and then?

OpenStudy (darkigloo):

\[-1=(-x+7)(2x+3)\]

OpenStudy (darkigloo):

now multiply out the right side

OpenStudy (anonymous):

u give me the answer and ill give u a medal

OpenStudy (darkigloo):

haha \[-1 = -2x ^{2} - 3x + 14x + 21\]

OpenStudy (darkigloo):

combine like terms

OpenStudy (darkigloo):

\[-1 = -2x ^{2} + 11x +21\] \[0 = -2x ^{2} + 11x + 22\]

OpenStudy (anonymous):

A 1/2 B 8 C -8 D 2

OpenStudy (darkigloo):

plug those for x and see which one equals 0

OpenStudy (anonymous):

answer plzzz

OpenStudy (darkigloo):

i dont have a calculator on me, and the calculator on the computer is annoying. if you really just want the answer, you could even just plug those x's into the equation youre given and see if they equal each other on both sides.